What is superphospate of lime and explain clearly abot it?

Grade:12th pass

10 Answers

KALYAN
8 years ago
 
Ca3(PO4)2 + 2H2SO4 = Ca(H2PO4)2 + 2CaSO4
2. a salt of phosphoric acid formed by incompletely replacing its acidic hydrogen atoms; acid phosphate; hydrogen phosphate
1. a mixture of the diacid calcium salt of orthophosphoric acid Ca(H2PO4)2 with calcium sulphate and small quantities of other phosphates: used as a fertilizer
 
SAI SARDAR
8 years ago
Dear Jyotheeswar,
           superphosphate or superphosphate of lime, Ca(H2PO4)2, is a compound produced by treating rock phosphate with sulfuric acid or phosphoric acid, or a mixture of the two.
Gaddam Chethan
8 years ago
Hello,
            superphosphate or superphosphate of lime, Ca(H2PO4)2, is a compound produced by treating rock phosphate with sulfuric acid or phosphoric acid, or a mixture of the two.
NARAYANA
8 years ago
my dear friend super phosphate of lime is also called calcium super phosphate and its formula Ca(H2PO4)2.
Prabhakar ch
8 years ago
 with calcium sulphate and small quantities of other phosphates: used as a fertilizer2)4PO21. a salt of phosphoric acid formed by incompletely replacing its acidic hydrogen atoms; acid phosphate; hydrogen phosphate2. a mixture of the diacid calcium salt of orthophosphoric acid Ca(H4+ 2CaSO 2)4PO2= Ca(H 4SO2+ 2H 2)4(PO3CaDear Jyotheeswar,
T.kumar
8 years ago
superphosphate or superphosphate of lime, Ca(H2PO4)2, is a compound produced by treating rock phosphate with sulfuric acid or phosphoric acid, or a mixture of the two. It is the principal carrier of phosphate, the form of phosphorus usable by plants, and is one of the world's most important fertilizers.
mohan
8 years ago
dear jyoti..
       superphospate  is a compound peoduced by treating rock with sulphuric acid or a mixture of the two.it is the principle carrier of phospate,the form of phosporous usable by plants,and is one of the worlds most important  fertilizers........
Gowri sankar
8 years ago
DEAR FRIEND, SUPERPHOSPATE IS A COMPOUND PRODUCED BY TREATING ROCK WITH SULPHURIC ACID.IT IS THE PRINCIPLE
raj
8 years ago
hellow jyothewar reddy superphosphate or superphosphate of lime, Ca(H2PO4)2, is a compound produced by treating rock phosphate with sulfuric acid or phosphoric acid, or a mixture of the two.
Y MANJUNADH
8 years ago
superphosphate of lime is also known as calcium superphosphate. this is one of the few water soluble phosphates.a mixture of calcium dihydrogen phospate is known as superphospate of lime.